7. find the slope of the line passing thru these points
a. (-2,4) and (-4,8)
b. (3,5) and (-5,-5)

Answer by Alan3354(69443)

7. find the slope of the line passing thru these points
a. (-2,4) and (-4,8)
slope m = diffy/diffx
m = (8-4)/(-4+2) = 4/(-2)
m = -2
-----------
b. (3,5) and (-5,-5)
Do it the same way.